A ball of mass 1kg is dropped from 20m height on ground and it rebounds to height 5m. Find magnitude of change in momentum during its collision with ground.(take g=10m/s2)
1
Expert's answer
2018-12-14T14:29:11-0500

If the ball is dropped from the height H, the speed near ground

v_1=√2gH=√(2×10 m/s^2 ×20 m)=20 m/s

The speed of the ball after rebound

v_2=√2gh=√(2×10 m/s^2 ×5 m)=10 m/s

So, the magnitude of change in ball’s momentum

Δp=mv_2+mv_1=1 kg×(10+20) m/s=30 (kg∙m)/s